OpenStudy (anonymous):

determine the common base of 4^2(x+5) - 11 = 245

OpenStudy (anonymous):

\[4^{2 (x + 5 )} - 11 = 245\]

OpenStudy (anonymous):

4^2(x+5) = 245 + 11 4^(2x + 10) = 256 4^(2x + 10) = 4^4 since this is an equality and bases are same i.e. 4 => powers have to be equal => 2x + 10 = 4 => 2x = 4 - 10 => x = -6/2 => x = -3

OpenStudy (anonymous):

thanks youuuuu!!

OpenStudy (anonymous):

u r welcome...☺

OpenStudy (anonymous):

but how come you used postive 11 its suppose to be negative.

OpenStudy (anonymous):

when u move it to the other side it becomes positive... the reason is that the exact process is 4^2(x+5) - 11 = 245 adding 11 to both sides so that -11 is eliminated from the LHS 4^2(x+5) - 11 + 11 = 245 + 11 4^2(x+ 5) = 245 + 11 hope it is clear now....